# The Effectiveness of Large Language Models in Transforming Unstructured Text to Standardized Formats

### v1

@@ 5.2 Limitations
While this study sheds light on the capabilities of Large Language Models (LLMs) in the converting recipes to Cooklang markdown, it is important to acknowledge the limitations of the study.
@@
    Model Versions: This study focuses on specific LLMs: GPT-4o, GPT-4o-mini, Llama:3.1:8b, and Llama:3.1:70b. To gain a more comprehensive understanding, further research should expand the scope to include a wider range of models.
@@
    Dataset Limitations: The study utilized a dataset of 32 distinct recipe samples from categories, which provides a foundation for initial analysis. However, to enhance the generalization of the results across different cuisines and recipe types, it would be beneficial to construct a larger, more comprehensive dataset. This expanded dataset could include a wider variety of recipes from diverse cultural backgrounds, cooking methods, and ingredient combinations. By evaluating the models on this larger dataset, we could gain more robust insights into their performance and better assess their applicability across a broader spectrum of culinary contexts.
@@
    Fine-tuning Potential: The study does not examine the potential for improvements resulting from task-specific fine-tuning, which could enhance performance and alter results. It would be beneficial for future research to include fine-tuning experiments for a more comprehensive evaluation. Fine-tuning large open-source models like Llama3.1 8b and 70b could potentially achieve performance comparable to GPT-4o or GPT-4o-mini, which highlights the importance of investigating this approach
Future research should address these limitations in order to gain a more comprehensive understanding of language model performance in recipe generation and related tasks.
@@ 6 Conclusion
Our research demonstrates the transformative potential of large language models (LLMs) in revolutionizing domain-specific content processing and generation. Through experimentation, we have shown that state-of-the-art models, particularly GPT-4o, can handle the intricate task of converting unstructured recipes into standardized Cooklang format with performance metrics—achieving a ROUGE-L score of 0.9722 and Word Error Rate of 0.0730 with few-shot prompting—showcasing not just incremental improvement but a quantum leap in automated recipe processing capabilities.
@@ 6.1 Future Research Directions
    Fine-tuning: The smaller models, like those in the Llama family, demonstrated difficulties in performing complex tasks such as accurate ingredient identification. However, the study suggests that targeted fine-tuning of these smaller models, such as Llama3.1:8b, could potentially narrow the performance gap in specific applications.
@@
    Broader model selection: While our current study established baseline performance using representative models (GPT-4o, GPT-4o-mini, Llama3.1:70b, and Llama3.1:8b), we plan to expand this analysis to include newer model iterations and additional open-source and proprietary models. This broader evaluation will help validate our findings and identify performance patterns across different model architectures and sizes.
@@
    Advanced Prompting Techniques: The exploration of alternative prompting techniques, such as multi-step prompting Fu et al. (2023) for recipe generation and conversion, and structured prompts that emulate the Cooklang format is a promising avenue for further research. That could improve every metric.
@@
    Dataset Expansion: The proposed improvements involve an expansion of the recipe dataset beyond the current 32 samples to encompass a more diverse range of cuisines, cooking methods, and ingredient combinations. This will enhance the generalization of the results. This approach could be also used to create a new dataset, it will convert recipe dataset into cooklang markdown. These kind of datasets could be used for finetuning large and small language models for this task. Furthermore, this approach facilitates the development of novel datasets, which contain the potential for the fine-tuning of both large and small language models, enabling more nuanced natural language understanding in culinary contexts. This transformation not only standardizes the recipe representation but also creates opportunities for cross-cultural culinary analysis and automated recipe processing applications.

### v2

@@ 5.3 Limitations
While this study sheds light on the capabilities of large language models (LLMs) in converting recipes to Cooklang markdown, it is important to acknowledge its limitations.
@@
    Model Versions: This study focuses on specific LLMs: GPT-4o, GPT-4o-mini, Llama:3.1:8b, and Llama:3.3:70b. To gain a more comprehensive understanding, further research should expand the scope to include a wider range of models. Specifically, testing the largest Llama3.1:405b model could provide insights into whether performance gaps between proprietary and open-source models narrow at the highest parameter counts. Our model selection was constrained by computational resources and aimed to provide a representative sample across both proprietary and open-source options rather than advocating for specific model families.
@@
    Dataset Limitations: The study utilized a dataset of 1098 distinct recipe samples from categories, which provides a foundation for initial analysis. However, to more enhance the generalization of the results across different cuisines and recipe types, it would be beneficial to construct a larger, more comprehensive dataset. This expanded dataset could include a wider variety of recipes from diverse cultural backgrounds, cooking methods, and ingredient combinations. By evaluating the models on this larger dataset, we could gain more robust insights into their performance and better assess their applicability across a broader spectrum of culinary contexts.
@@
    Fine-tuning Potential: The study does not examine the potential for improvements resulting from task-specific fine-tuning, which could enhance performance and alter results. It would be beneficial for future research to include fine-tuning experiments for a more comprehensive evaluation. Finetuning Large Language models for recipe generation from the study by Vij et al. Vij et al. (2025) demonstrates the significant impact of task-specific fine-tuning on model performance for specialized domains like recipe generation. While fine-tuning generally improved metrics like step complexity across models of various sizes (from SmolLM-135M to Phi-2), their research revealed nuanced trade-offs in performance metrics, similar as ours. Their comparative analysis across models demonstrated that smaller models can sometimes achieve comparable performance to larger ones after fine-tuning, indicating that model size alone doesn’t determine fine-tuning efficacy, which supports our claims that finetuning smaller models could lead to similar results as models from the OpenAI family.
Future research should address these limitations to gain a more comprehensive understanding of language model performance in recipe generation and related tasks.
@@ 6 Conclusion
Our research demonstrates the transformative potential of large language models (LLMs) in revolutionizing domain-specific content processing and generation. Through experimentation, we have shown that state-of-the-art models, particularly GPT-4o, can handle the intricate task of converting unstructured recipes into standardized Cooklang format with performance metrics—achieving a ROUGE-L score of 0.8128 and Word Error Rate of 1.5255 with few-shot prompting—showcasing not just incremental improvement but a quantum leap in automated recipe processing capabilities.
@@ 6.1 Future Research Directions
    Fine-tuning: The smaller models, like those in the Llama family, demonstrated potential in performing complex tasks such as accurate ingredient identification. However, the study suggests that targeted fine-tuning of these smaller models, such as Llama3.1:8b, could potentially narrow the performance gap in specific applications.
@@
    Broader model selection: While our current study established baseline performance using representative models (GPT-4o, GPT-4o-mini, LLama3.3:70b, and Llama3.1:8b), we plan to expand this analysis to include newer model iterations and additional open-source and proprietary models. This broader evaluation will help validate our findings and identify performance patterns across different model architectures and sizes.
@@
    Advanced Prompting Techniques: The exploration of alternative prompting techniques, such as multi-step prompting Fu et al. (2023) for recipe generation and conversion, and structured prompts that emulate the Cooklang format, is a promising avenue for further research. That could improve every metric.
@@
    Dataset Expansion: The proposed improvements involve an expansion of the recipe dataset beyond the current 1098 samples to encompass a more diverse range of cuisines, cooking methods, and ingredient combinations. This approach could be also used to create a new dataset. It will convert the recipe dataset into a cooklang markdown. These kinds of datasets could be used for finetuning large and small language models for this task. Furthermore, this approach facilitates the development of novel datasets, which contain the potential for the fine-tuning of both large and small language models, enabling more nuanced natural language understanding in culinary contexts. This transformation not only standardizes the recipe representation but also creates opportunities for cross-cultural culinary analysis and automated recipe processing applications.
@@ 6.2 Closing Remarks
7 Acknowledgment
The authors gratefully acknowledge Alexey Dubovskoy (dubadub) and CapnDan for providing their personal cooklang recipes, which served as the evaluation backbone for this research. We also extend our sincere appreciation to the broader cooklang community for their ongoing contributions and commitment to structured recipe formatting. The diverse collection of recipes made available by this community has been instrumental in validating our methodological approach and enhancing the robustness of our findings.